The price of a 2-hour plumbing repair job is $130 and the price of a 5-hour job is $250. Find and interpret the slope of the line through (2, 130) and (5, 250).
Answer by checkley77(12844)

SLOPE=(250-130)/(5-2)=120/3=$40 PER HOUR IS THE PLUMBING RATE.
2*40=80
130-80=50
5*40=200
250-200=50
THEREFORE THERE MUST BE A MINIMUM CALL RATE OF $50, IN ADDITION TO THE $40 PER HOUR RATE.